Calculate the change in entropy when 12g of water at -10oC is raised to 100oC and convert to steam at 100oC. Assume that specific heat of water cwater = 4190 J·g-1·K-1, heat of vaporization Lv = 2256 kJ/kg, specific heat of ice cice = 2220 J·g-1·K-1, heat of fusion LF = 334 kJ/kg.
1
Expert's answer
2019-06-19T10:30:18-0400

The total entropy change is:


"Q = m * (c_i*
\Delta T_1 + H_f + c_w*\Delta T_2 + H_v)"

"Q = 12 * (2220*10 + 334 + 4190*100 + 2256)"

"Q = 5325 (kJ)"